# Planet formation regulated by galactic-scale interstellar turbulence

### v2

@@ 3.2 Observational comparison
We wish to compare our model to observed discs with comparable stellar ages. The ages of star forming regions can both be observationally uncertain (e.g. Soderblom et al., 2014) and should exhibit physical scatter given the finite duration of the star formation process. However, nominally ages for the youngest populations of stars in the $\rho$ Ophiuchus complex are $\sim 0.1-1$ Myr (Luhman & Rieke, 1999) in Taurus are $\sim 1{-}3$ Myr (Krolikowski et al., 2021), in Lupus are $\lesssim 3$ Myr (Galli et al., 2020) in Chameleon I have been estimated at $\sim 1{-}2$ Myr (Galli et al., 2021b) and $\sim 3{-}6$ Myr (Luhman, 2007), and Upper Sco are $\sim 5-10$ Myr (Pecaut & Mamajek, 2016; Luhman & Esplin, 2020). We compare these regions to our model at ages, $0.5$, $1$, $2$, $3$ and $5$ Myr respectively; these choices should be considered illustrative, not precise. For total disc masses we take $M_{\mathrm{disc}}=10^{2}M_{\mathrm{dust}}$, where $M_{\mathrm{dust}}$ is as quoted by Manara et al. (2023), inferred from the mm continuum flux with standard assumptions for opacity and temperature. In order to draw comparisons with each, we perform a Gaussian kernel density estimate with Scott’s rule for bandwidth selection (Scott, 1992). Where we compare to data directly, we weight the model data points to yield the same distribution in the stellar mass axis as for the observational dataset. In practice, this means dividing the stellar mass distribution into bins, and defining the weight of model data points in each bin $i$ to be $N_{\mathrm{obs},i}/N_{\mathrm{mod},i}$, where $N_{\mathrm{obs},i}$ and $N_{\mathrm{model},i}$ are the number of observed and model points in bin $i$ respectively. The results of these comparisons are shown in Figure 3, in which we find remarkable agreement with the observed disc mass, stellar accretion rate, and outer disc radius distributions. We explore these findings in greater detail as follows.
@@ 3.2.2 Stellar accretion rates
The accretion rates from our model are entirely determined by the disc mass distribution, since we have imposed $\dot{M}_{\mathrm{acc}}\propto M_{\mathrm{disc}}/\tau_{\mathrm{acc}}$. They are therefore not an independent test of our model, but a validation or repudiation of this assumption. Given that we have anchored $\tau_{\mathrm{acc}}$ to observations, and we have found observational agreement with the disc mass distribution, as expected we also find for the majority of regions we have excellent agreement for accretion rates (middle row, Figure 3). We note minor tension particularly in the case of Chameleon I, for which some accretion rates are somewhat larger than predicted by our model. This is plausibly the result of a younger population than the comparison age we have adopted. For example, Galli et al. (2021b) argue that the older esimate obtained by Luhman (2007) may be explained by systematic distance underestimate (then Chameleon I may be $\sim 1{-2}$ Myr old). Overall, we find excellent agreement between stellar accretion rates in our model and the observed samples. Since our model does not contain any stellar mass dependent disc physics (except the BHL accretion rate), this makes a strong argument in favour of BHL accretion as the origin of disc assembly.
@@ 3.2.3 Outer disc radii
We can also consider the distribution of gaseous disc radii $R_{\mathrm{disc}}$ we infer from our model. However, the appropriate definition of $R_{\mathrm{disc}}$ is less clear than for disc masses or stellar accretion rates in terms of the correspondence between the model and observational data. The observed outer radius in dust is dependent on variations in the dust opacity (Rosotti et al., 2019) and in gas molecular tracers, such as CO, outer radii are dependent on thermochemistry (Trapman et al., 2023). Both are dependent on the definition of ‘outer radius’, often defined by the radius that encloses some fraction of the total flux. In our simple models, we do not explicitly define an outer radius. A sensible estimate for $R_{\mathrm{disc}}$ is the accretion radius $R_{\mathrm{acc}}$ at the most recent time for which the disc is accumulating mass (i.e. $\dot{M}_{\mathrm{BHL}}>\dot{M}_{\mathrm{acc}}$), and this is the definition we adopt in this work and in constructing Figure 3. In the absence of other disc evolution processes, such as viscous spreading (Lynden-Bell & Pringle, 1974), we actually expect $R_{\mathrm{acc}}>R_{\mathrm{disc}}$, since the outer radius depends on the fraction of residual angular momentum from the accreted ISM. With these caveats, we proceed with the assumption $R_{\mathrm{acc}}\approx R_{\mathrm{disc}}$ in our comparison with observations in the main article.
@@ 3.2.4 Disc lifetimes
For a fair comparison of disc property distributions, we must also confirm that the total fraction of discs included in our sample is similar in our model to those in star forming regions of a given age. We therefore show the disc fraction as a function of time in Figure 6, by the definition $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$. We confirm that our model yields reasonable disc fractions over the range of times we consider in our disc property comparisons. We find a very similar disc fraction evolution to the exponential decay relationships suggested in the literature, typically with dispersal time-scale $\tau_{\mathrm{disc}}\sim 3{-}5$ Myr (Haisch et al., 2001; Ribas et al., 2015).
Figure 6: Fraction of surviving discs in our model (black line), defined to be discs which have a mass $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$ for stars with masses in the range $0.1{-}3\,M_{\odot}$. We show the evolution of disc fraction given by $\exp\left(-t/\tau_{\mathrm{disc}}\right)$ for $\tau_{\mathrm{disc}}=3$ Myr (Haisch et al., 2001) and $5$ Myr (Ribas et al., 2015) as blue and red lines respectively.
Interestingly, we find that the initial disc fraction is not $100$ percent in our model, but is maximal at $\sim 90$ percent. Observations indicate that, even for very young disc populations, the disc fraction never exceeds $\sim 80$ percent (Michel et al., 2021). However, when accounting for close binaries, this fraction could be closer to $100$ percent for single stars (Kraus et al., 2012). The model prediction of a somewhat suppressed initial disc fraction should therefore not be interpreted as a unique signature of BHL accretion, and may also depend on the presence or absence of a primordial disc. We also find a somewhat flatter disc fraction as a function of time than a simple exponential decay with linear time exponent. Some authors to suggest such a flattened functional form for the disc fraction (Pfalzner et al., 2022), supported by a non-negligible disc fractions in clusters that are tens of Myr old (Galli et al., 2021a). Such surviving discs can be easily understood in the context of the stochastic accretion process that defines our model.
@@
3.3 Replenished mass fraction:
As discussed in the main article, it is not clear immediately from Figure 3 whether disc properties are set by a continuous process of BHL accretion. Does BHL just set the initial disc properties that then simply evolve over our imposed time-scale $\tau_{\mathrm{acc}}$, or is disc material constantly replenished over the disc lifetime? This is an important distinction because if disc material is constantly being replenished then, even with well-determined initial conditions, the distribution of disc properties can never be a direct probe of isolated disc evolution.
@@ 3.3 Replenished mass fraction:
Figure 7: The fraction of mass that has been accreted in the last half of the disc life-time, $f_{M,1/2}$, as a function of time for $500$ discs in our model. On the left, we show the time evolution of $f_{M,1/2}$ for a random subset of $100$ discs for clarity. These lines are coloured by the ratio of the disc mass to stellar mass $M_{\mathrm{disc}}/m_{*}$, shown in the colour bar along the top. On the right hand side we show a cumulative fraction for $f_{M,1/2}$ at $1$, $2$, $3$ and $5$ Myr. Only discs with a total mass $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$ are included in this cumulative fraction, lower mass discs are assumed to be dispersed.
@@
Applying this metric in Figure 7, we find a substantial fraction of discs composed of a high proportion of recently captured ISM material. We therefore conclude that disc properties are constantly ‘reset’. This would suggest that the distributions of observed disc properties should not be used to test of theoretical models for isolated disc evolution.
3.4 Sensitivity to the initial disc mass:
@@
In drawing the above conclusions, we have made the assumption that the initial disc mass is negligible compared to the subsequently accreted material. As discussed, this assumption is justified in the desire for a ‘clean’ test of the disc properties resulting from BHL accretion, retaining a simple model with few free parameters, and by the coincidence that an alternative would imply. However, given that this assumption is also unconventional, we can ask whether our conclusion regarding the degree of disc replenishment holds in the event of the coincidence that the primordial disc mass is comparable to the subsequently captured mass.
To investigate the sensitivity of our conclusions on disc replenishment to this assumption, we repeat our calculations with disc masses initially similar to the observed distribution in young star forming regions. In practice, we adopt a log-normal distribution for the disc mass with $1$ dex of scatter around the relationship
$M_{\mathrm{disc}}(t=0)=0.01\left(\frac{m_{*}}{1\,M_{\odot}}\right)^{2}\,M_{\odot}.$ (11)
We show the resultant disc properties in  8, which again are in good agreement with the observed disc properties. Our adopted initial disc masses are approximately the maximum we could adopt while still remaining consistent with the observed distribution of disc masses. The outcome for the half-life disc replenishment factor is shown in Figure 9. Although in this case $f_{\mathrm{M,1/2}}$ is somewhat reduced early on in the disc evolution, overall differences in the distribution over time is minor. The fraction of discs with at least some recently accreted material remains similar, while we find that $f_{\mathrm{M,1/2}}\sim 0.2{-}0.45$ throughout the disc lifetime. Therefore, even under the assumption that the primordial disc is coincidentally similar in mass to the subsequently accreted one, and ignoring all internal depletion processes, we still expect a substantial fraction of discs to be composed of recently captured disc material. Our conclusion that the present day disc properties are not a good probe of internal disc physics therefore appears robust.
Figure 10: The column number density of material assuming a mean molecular mass $\mu=2.3$ and $\epsilon_{\mathrm{geo}}=1$ following equation 21 for the surviving discs in our fiducial model. The colours of the lines represent the age of the discs. The red vertical dashed line represents an approximate column density for photodissociation of 12CO, based on a self-shielding column of $10^{15}$ cm-2 (van Dishoeck & Black, 1988) and abundance relative to hydrogen of $10^{-5}$ (Trapman et al., 2023)
3.5 Disc turbulence and angular momentum transport:
@@
In our model, we have assumed that disc mass is accreted onto the star with a fixed time-scale $\tau_{\mathrm{acc}}$ via an undefined angular momentum transport process. This process could be internal to the star-disc system, driven by magnetic winds or due to (magneto-)hydrodynamic instability (Tabone et al., 2022; Lesur et al., 2023). However, the ISM itself could also be the origin of protoplanetary disc turbulence. The degree to which angular momentum transport is mediated by turbulence remains an open question, but measurements indicate it could be the origin of stellar accretion in at least some discs (Rosotti, 2023).

### v3

@@ 3.2 Observational comparison
We wish to compare our model to observed discs with comparable stellar ages. The ages of star forming regions can both be observationally uncertain (e.g. Soderblom et al., 2014) and should exhibit physical scatter given the finite duration of the star formation process. However, nominally ages for the youngest populations of stars in the $\rho$ Ophiuchus complex are $\sim 0.1-1$ Myr (Luhman & Rieke, 1999) in Taurus are $\sim 1{-}3$ Myr (Krolikowski et al., 2021), in Lupus are $\lesssim 3$ Myr (Galli et al., 2020) in Chameleon I have been estimated at $\sim 1{-}2$ Myr (Galli et al., 2021b) and $\sim 3{-}6$ Myr (Luhman, 2007), and Upper Sco are $\sim 5-10$ Myr (Pecaut & Mamajek, 2016; Luhman & Esplin, 2020). We compare these regions to our model for ages of the individual stars (not the regions), $0.5$, $1$, $2$, $3$ and $5$ Myr respectively; these choices should be considered illustrative, not precise. For total disc masses we take $M_{\mathrm{disc}}=10^{2}M_{\mathrm{dust}}$, where $M_{\mathrm{dust}}$ is as quoted by Manara et al. (2023), inferred from the mm continuum flux with standard assumptions for opacity and temperature. In order to draw comparisons with each, we perform a Gaussian kernel density estimate with Scott’s rule for bandwidth selection (Scott, 1992). Where we compare to data directly, we weight the model data points to yield the same distribution in the stellar mass axis as for the observational dataset. In practice, this means dividing the stellar mass distribution into bins, and defining the weight of model data points in each bin $i$ to be $N_{\mathrm{obs},i}/N_{\mathrm{mod},i}$, where $N_{\mathrm{obs},i}$ and $N_{\mathrm{model},i}$ are the number of observed and model points in bin $i$ respectively. The results of these comparisons are shown in Figure 3, in which we find remarkable agreement with the observed disc mass, stellar accretion rate, and outer disc radius distributions. We explore these findings in greater detail as follows.
@@ 3.2.2 Stellar accretion rates
The accretion rates from our model are entirely determined by the disc mass distribution, since we have imposed $\dot{M}_{\mathrm{acc}}\propto M_{\mathrm{disc}}/\tau_{\mathrm{acc}}$. They are therefore not an independent test of our model, but a validation or repudiation of this assumption. Given that we have anchored $\tau_{\mathrm{acc}}$ to observations, and we have found observational agreement with the disc mass distribution, as expected we also find excellent agreement for accretion rates for the majority of regions (middle row, Figure 3). We note minor tension particularly in the case of Chameleon I, for which some accretion rates are somewhat larger than predicted by our model. This is plausibly the result of a younger population than the comparison age we have adopted. For example, Galli et al. (2021b) argue that the older esimate obtained by Luhman (2007) may be explained by systematic distance underestimate (then Chameleon I may be $\sim 1{-2}$ Myr old). Overall, we find excellent agreement between stellar accretion rates in our model and the observed samples.
@@ 3.2.3 Outer disc radii
We can also consider the distribution of gaseous disc radii $R_{\mathrm{disc}}$ we infer from our model. However, the appropriate definition of $R_{\mathrm{disc}}$ is less clear than for disc masses or stellar accretion rates in terms of the correspondence between the model and observational data. The observed outer radius in dust is dependent on variations in the dust opacity (Rosotti et al., 2019) and in gas molecular tracers, such as CO, outer radii are dependent on thermochemistry (Trapman et al., 2023). Both are dependent on the definition of ‘outer radius’, often defined by the radius that encloses some fraction of the total flux. In our simple models, we do not explicitly define an outer radius. Simulations (Krumholz et al., 2006; Pelkonen et al., 2024) and analytic estimates (Padoan et al., 2024) indicate that $R_{\mathrm{disc}}\sim R_{\mathrm{acc}}\propto m_{*}$ when this disc is accreted from a turbulent medium, with the disc radius typically a factor order unity smaller than the accretion radius. With the above caveats, we therefore attempt to compare $R_{\mathrm{acc}}$ with the observed outer disc radii. While there are fewer measurements of gaseous disc outer radii in local star forming regions, in the bottom row of Figure 3 we compare our model to the empirically-inferred relationship for the outer gas disc radius as inferred from CO molecular line measurements: $R_{\mathrm{CO}}=250\left(m_{*}/1\,M_{\odot}\right)^{0.9}$ au (Andrews, 2020) typically a factor $\sim 2{-}3$ larger than the dust component, although with considerable scatter and some examples of much more extended CO radii (Sanchis et al., 2021). We again find a remarkable similarity between our model results and the observed disc properties, with a tail of large $R_{\mathrm{acc}}>R_{\mathrm{CO}}$. The agreement between the normalisation in our model and the data may be somewhat coincidental, since 12CO is usually optically thick so that $R_{\mathrm{CO}}$ probably overestimates where the bulk of the mass is, while $R_{\mathrm{acc}}$ also somewhat overestimates $R_{\mathrm{disc}}$. It is nonetheless clear that both the scaling and approximate normalisation of the $m_{*}-R_{\mathrm{disc}}$ relationship are predicted directly by BHL accretion.
@@ 3.2.4 Disc lifetimes
For a fair comparison of disc property distributions, we must also confirm that the total fraction of discs included in our sample is similar in our model to those in star forming regions of a given age. We therefore show the disc fraction as a function of time in Figure 6, by the definition $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$. We confirm that our model yields reasonable disc fractions over the range of times we consider in our disc property comparisons. We find a very similar disc fraction evolution to the exponential decay relationships suggested in the literature, typically with dispersal time-scale $\tau_{\mathrm{disc}}\sim 3{-}5$ Myr (Haisch et al., 2001; Ribas et al., 2015). In Figure 6, we also show the disc fractions we obtain when multiplying the threshold mass by $1/3$ and $3$, finding only a moderate difference of $\sim 10$ percent between the most extreme choices.
Figure 6: Fraction of surviving discs in our model (solid black line), defined to be discs which have a mass $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$ for stars with masses in the range $0.1{-}3\,M_{\odot}$. We also vary this threshold, shown by the dashed and dotted lines (factor $1/3$ and $3$ respectively). We show the evolution of disc fraction given by $\exp\left(-t/\tau_{\mathrm{disc}}\right)$ for $\tau_{\mathrm{disc}}=3$ Myr (Haisch et al., 2001) and $5$ Myr (Ribas et al., 2015) as blue and red lines respectively.
Interestingly, we find that the disc fraction never reaches $100$ percent in our model, but is maximal at $\sim 90$ percent. Observations indicate that, even for very young disc populations, the disc fraction never exceeds $\sim 80$ percent (Michel et al., 2021). However, when accounting for close binaries, this fraction could be closer to $100$ percent for single stars (Kraus et al., 2012). The model prediction of a somewhat suppressed early disc fraction should therefore not be interpreted as a unique signature of BHL accretion, and may also depend on the presence or absence of a primordial disc. We also find a somewhat flatter disc fraction as a function of time than a simple exponential decay with linear time exponent. Some authors to suggest such a flattened functional form for the disc fraction (Pfalzner et al., 2022), supported by a non-negligible disc fractions in clusters that are tens of Myr old (Galli et al., 2021a). Such surviving discs can be easily understood in the context of the stochastic accretion process that defines our model.
@@
3.3 Replenished mass fraction
It is not clear immediately from Figure 3 whether disc properties are set by a continuous process of BHL accretion. Does BHL just set the initial disc properties that then simply evolve over our imposed time-scale $\tau_{\mathrm{acc}}$, or is disc material constantly replenished over the disc lifetime? This is an important distinction because if disc material is constantly being replenished then, even with well-determined initial conditions, the distribution of disc properties can never be a direct probe of isolated disc evolution.
@@ 3.3 Replenished mass fraction
Figure 7: The fraction of mass that has been accreted in the last half of the disc life-time, $f_{M,1/2}$, as a function of time for $500$ discs in our model. On the left, we show the time evolution of $f_{M,1/2}$ for all $500$ discs (faint lines), with a random subset of $10$ discs shown boldly for clarity. These lines are coloured by the ratio of the disc mass to stellar mass $M_{\mathrm{disc}}/m_{*}$, shown in the colour bar along the top. On the right hand side we show a cumulative fraction for $f_{M,1/2}$ at $1$, $2$, $3$ and $5$ Myr. Only discs with a total mass $M_{\mathrm{disc}}>3\times 10^{-5}\,M_{\odot}$ are included in this cumulative fraction, lower mass discs are assumed to be dispersed.
Applying this metric in Figure 7, we find a substantial fraction of discs composed of a high proportion of recently captured ISM material. On the left panel we show how the replenishment fraction evolves over time for a random subset of discs. Typically, $f_{\mathrm{M,1/2}}\approx 0.5$ early on, then the fraction drops once the phase in which stars occupy the bound star forming region finishes. As the disc reduces in mass, discs often see an increase in $f_{\mathrm{M,1/2}}$ later in their lifetime, when relatively modest BHL accretion rates can still provide substantial replenished mass compared to the existing disc mass. Not all discs follow this pattern however, with some discs having both high replenishment fractions ($\gtrsim 0.5$) and considerable disc masses ($M_{\mathrm{disc}}/m_{*}\gtrsim 10^{-3}$) at relatively old ages ($\gtrsim 5$ Myr). The right panel shows that the fraction of discs composed of mostly replenished material drops from $\sim 70$ percent at 1 Myr, to $\sim 30$ percent at $2$ Myr and $\sim 20$ percent for discs older than $3$ Myr. This fraction therefore remains substantial throughout the disc lifetime.
@@
We conclude that disc properties are constantly ‘reset’ for many discs in local star forming regions during their first few Myr of evolution. This would suggest that the distributions of observed disc properties should not be used as a test of theoretical models for isolated disc evolution.
3.4 Caveats and tests
3.4.1 Class 0 phase and the initial disc mass
@@
To simplify our model, we have assumed that a star initially forms instantaneously, without a period of growth that corresponds to the protostellar and class 0 phase. Indeed, for our fiducial model, we have made the assumption that the initial disc mass is negligible compared to the subsequently accreted material. As discussed, this assumption is justified in the desire for a ‘clean’ test of the disc properties resulting from BHL accretion, retaining a simple model with few free parameters, and by the coincidence that an alternative would imply. However, given that this assumption is also unconventional, we can ask whether our conclusion regarding the degree of disc replenishment over time holds in the event of the coincidence that the primordial disc mass is comparable to the subsequently captured mass.
We explore this assumption in Appendix D, repeating our experiment but assuming massive initial discs. While we still do not capture the class 0 disc assembly phase, this is equivalent to the opposite assumption as for our fiducial model; that the majority of the young discs mass is assembled within the first $\sim 10^{5}$ yr. From Figure 15, we conclude that the replenishment fraction is only influenced for young discs ($\lesssim 1$ Myr), after which the distributions remain similar (cf. Figure 7). Our conclusion is that disc replenishment remains an important process throughout the disc lifetime therefore does not depend on this assumption.
3.4.2 Internal winds
As discussed in Section 2.6, photoevaporative winds driven by high energy photons from the accreting star may suppress the BHL accretion flow (e.g. Edgar & Clarke, 2004). While quantifying the degree to which BHL accretion is suppressed is a complex problem, and should be the subject of a dedicated study, in Appendix E we outline a simplified approach.
We discuss the results of including the internally driven wind in detail in Appendix E, but highlight two key conclusions as follows. The first is that photoevaporation leads to more discrete, episodic periods of accretion, particularly for high mass stars (Figure 18). When stars are in a sufficiently low density environment, the wind completely suppresses accretion in our model, although in reality some accretion may still be possible. Secondly, the fraction of the disc mass that has been recently accreted is reduced by including the wind at late times (Figure 21 compared to Figure 7). Despite including this wind, our estimate suggests that $\sim 20$ percent of discs are still composed of mostly replenished material at an age of $2-3$ Myr. This fraction is not very strongly influenced because the periods where the majority of the mass is accumulated are the rapid BHL accretion phases where the dense accretion column is not disrupted by the wind. Nonetheless, clearly it is plausible that internally driven winds can disrupt the BHL accretion flow, at least for stars $\gtrsim 2\,M_{\odot}$ (Figure 17), and therefore this represents an important topic for future study.
3.4.3 External stellar feedback
Apart from internal irradiation, we have excluded heating of the ambient medium by neighbouring stars. For older regions, with a lower ambient density, and more massive regions, with a larger number of OB stars, feedback should eventually become important for setting the ambient gas density and sound speed. This will in turn suppress BHL accretion, perhaps in conjunction with the internal wind discussed in Section 3.4.2. Specifically, for feedback-heated regions, the sound speed and relative gas velocity would be enhanced, reducing $R_{\mathrm{BHL}}$ and possibly changing the mode of accretion from BHL to disc sweep-up (Section G.3). While we do not expect this to strongly influence our results for the local regions which are the focus of this work (although perhaps more so in Upper Sco), this is clearly a factor that must be considered when interpreting disc properties in different environments (e.g. van Terwisga & Hacar, 2023).
3.4.4 Lower limit of star forming region mass
As shown in Figure 11, star forming regions of lower mass collapse over a shorter time-scale, with higher density. We might then expect the mass accumulation for these regions to be more dominated by this early phase. We have chosen a minimum mass of a collapsing GMC to be $10\,M_{\odot}$, but if we reduce this we include more stars in lower mass regions that may then be more dominated by the initial stages of BHL accretion. However, we test the sensitivity of our results to this lower mass threshold in Appendix F, finding that in a statistical sense our results are not strongly sensitive to our choice.
3.4.5 Other caveats for the BHL accretion rate
There remain numerous caveats for our assumed BHL accretion rate. For example, complicating factors include magnetic fields, self-gravity of the gas, and a change in accretion mechanism when the BHL radius is within the disc radius (sweep-up). We discuss these issues in greater detail in Appendix G; generally we expect deviations of order unity from the nominal BHL accretion rate.
3.5 Results summary
The qualitative results of our disc evolution calculations can be summarised concisely: BHL accretion is capable of reproducing observed disc masses, accretion rates, radii and lifetimes as a function of stellar mass and time. The BHL accretion rate is sufficient to have recently re-supplied a considerable fraction ($\sim 20{-}70$ percent) of discs of age $1{-}3$ Myr with a high fraction ($\gtrsim 50$ percent) of replenished material.
4 Discussion
4.1 Disc turbulence and stellar accretion
@@
In our model, we have assumed that disc mass is accreted onto the star with a fixed time-scale $\tau_{\mathrm{acc}}$ via an undefined angular momentum transport process. This process could be internal to the star-disc system, driven by magnetic winds or by viscosity generated by (magneto-)hydrodynamic instability (Tabone et al., 2022; Lesur et al., 2023). Here we assess whether BHL accretion, and late stage in-fall more generally, could drive turbulence and/or stellar accretion in protoplanetary discs. In terms of the latter, the case against BHL accretion as the origin for stellar accretion (i.e. the accretion of the disc material onto the star) was concisely and influentially laid out by Hartmann et al. (2006), based on two arguments. We consider both of these arguments in Sections 4.1.1 and 4.1.2.
4.1.1 Angular momentum transport to the inner disc
The first argument against BHL accretion as a driver of stellar accretion is that accreted material would presumably need to have negligible angular momentum in order to accrete close to the stellar radius. Both simulations of BHL accretion in a turbulent medium (Krumholz et al., 2006; Kuffmeier et al., 2023; Pelkonen et al., 2024) and the fact that disc radii far exceed the stellar radius imply that this is not the case. This would appear to demonstrate that BHL accretion cannot set stellar accretion rates. However, observationally it is clear that, at least for the highest accretors, (magnetically-driven) winds can be launched from radii $\lesssim 10$ au (see Section 4.1 of Pascucci et al., 2023), and other internal processes may also drive turbulence in the inner disc (e.g. Carr et al., 2004; Najita et al., 2009; Ilee et al., 2014; Bosman et al., 2023). Then in-falling material only needs to reach down to these radii in order to eventually accrete onto the host star. Even if the initial angular momentum of material is significantly larger than this, some angular momentum transport in the outer disc may be driven by in-fall in two ways.
The first way is the misalignment of the angular momentum vector of material accreted at different times. For example, Kuffmeier et al. (2024) recently showed that the angular momentum vector of accreted material in a turbulent medium is typically decorrelated on time-scales $\lesssim 1$ Myr. Clearly, adding fluid elements together with the same magnitude of angular momentum but offset angular momentum vectors will reduce the total specific angular momentum of the system. This can simply be understood as a delayed form of BHL directly onto the star. Such a process would presumably lead to shrinking of the disc over time, and may therefore leave a similar imprint in disc populations as MHD driven winds (Tabone et al., 2022; Manara et al., 2023). It would also result in misalignment between inner and outer discs, which is indeed evident in many discs (Bohn et al., 2022; Villenave et al., 2024). Whether or not this results in a continuous or bursty supply of material to the inner disc would depend on whether warps instigated by in-falling material decay on comparatively long time-scales (e.g. the $\sim 10^{5}$ years estimated by Kimmig & Dullemond, 2024), or whether associated instabilities drive high levels of turbulence that rapidly decays (Deng et al., 2020; Fairbairn & Ogilvie, 2023).